If angles A and B are complimentary what are the measures of angles A and B if angle A is three times the measure of angle B.
Answer by nerdybill(7384)

If angles A and B are complimentary what are the measures of angles A and B if angle A is three times the measure of angle B.
.
Let B = measure of angle B
then
3B = measure of angle A
.
B + 3B = 90
4B = 90
B = 90/4
B = 22.5 degrees
.
Angle A:
3B = 3(22.5) = 67.5 degrees